# Quantum entanglement, coherence and Dark Matter

### v1

@@ II.1 The two-atom problem
with $\theta$ and $\theta^{\prime}$ denoting the orientations of the first and second absorbed photons, respectively, relative to the $z$-axis (the absorption of the second photon is also assumed to stem from a one-photon field). Here again, there is angular correlation, but this time between successively absorbed photons. Superabsorption will also play an important role when our model is applied to more general considerations in Sec. IV.
II.2 Generalization to $n$ atoms
For the two-atom problem only two permutation symmetries are available for the Dicke states: the totally symmetric species is shared by the triplet $r=1$ (equations 7-9), while the lone $r=0$ singlet is anti-symmetric (equation 10). In the more general case where the gas is composed of an arbitrary number of atoms $n$ a larger number of symmetries, tied to the cooperative number $r$, exist for the Dicke states. For example, it can be shown that a four-atom system would allow for five non-degenerate totally symmetric $r=2$ states, three levels of triply-degenerate $r=1$ states, and one level of doubly-degenerate $r=0$ dark states.
The general case for an arbitrary number of atoms $n$ is shown in Figure 1 along with the expected transition rates in the small sample (i.e., $kz_{0}\rightarrow 0$) or coherent limits (taken from Dicke’s original paper [1]). As can be seen in the figure, it is generally observed that, for a given $m$ value, transitions are allowed only between states sharing the same cooperation number $r$, while transition rates systematically decrease as one moves rightward to lower $r$ values. This behavior is consistent with that observed for the two-atom problem treated above.
Figure 1: The coherent radiation rates between Dicke states in the small sample (i.e., when $kr\rightarrow 0$) or coherent limits. The number of atom contained in the radiating gas is $n$, while the states within the different coherent cascades are identified through the cooperative number $20\leq r\leq n/$ 2 and the inversion number $-r\leq m\leq r$. The radiative intensity (or rate) of a single atom is denoted by $I_{0}$. Taken from Dicke’s original 1954 paper [1].
Although it is in principle possible to extend the spontaneous emission rates analysis for finite separations previously performed for the two-atom case (in Sec. II.1) to the four- or $n$-atom problems using appropriate Dicke states, the calculations quickly become prohibitive even for a restricted number of radiators. Such analysis would, for example, allow us to estimate the amount of energy trapped in the $n$-atom gas when starting with some initial conditions (e.g., equally populating the $m=0$ states) and establish any potential tendency for the radiation intensity as $n$ increases for a given spatial distribution.
It is, however, possible to determine the ratio of numbers of photons emitted in the small and infinite-size samples when $m=0$, as was done in equation (23) for the two-atom problem. For this, we use the known level of degeneracy of states as a function of the cooperative number $r$ [1]
@@ II.2 Generalization to $n$ atoms
with the fact that, for $n$ even, the number of photons emitted from $m=0$ for the small and infinite-size samples are $r$ and $n/2$, respectively. We thus find for the energy trapping efficiency 222for $n$ odd, we have for the energy trapping efficiency $\eta=1-\frac{2}{n-1}\left[\sum_{r=1/2}^{n/2}g_{r}\left(r-1/2\right)\right]/\left(\sum_{r=1/2}^{n/2}g_{r}\right)$ from $m=-1/2$.
@@
One can readily verify that $\eta=1/2$ for the two-atom case, while it is $0.69$ for $n=10$, $0.88$ for $n=100$, etc. It is therefore clear that the combination of coherence and the entanglement of a large number of atoms can potentially lead to a significant amount of energy trapping within the gas. We also note that this effect gets more pronounced as the system starts at lower $m$ values. For example, we find $\eta=1-1/n$ for equal initial occupation probabilities of the $m=-n/2+1$ states [1].
But one of our goals is to assess the level of energy trapping within the gas for finite and more realistic atomic separations in the interstellar medium, i.e., beyond the small sample limit as we have done so far. It is already obvious through our study of the two-atom system that the general case where $kz_{0}\neq 0$ will lead to a non-zero transition rate from the |$\Ket{0,0}_{\theta}$ ‘dark’ state (see equations 10 and 16). This has for implication that all the energy initially stored in the system will eventually leak away. That is, $\eta=0$ in the steady-state and no energy trapping is realized in the long run.
III Equilibrium conditions
To further our study of trapping and leaking of internal energy we move to density matrix framework. More precisely, we will make use of the master equation often used within the context of superradiance [20, 24]
$\frac{d\hat{\rho}}{dt}$ $=\frac{1}{i\hbar}\hbar\omega_{0}\sum_{j}\left[\hat{R}^{3}_{j},\hat{\rho}\right]-\frac{1}{i\hbar}\sum_{i\neq j}\Omega_{ij}\left(k\Delta r_{ij}\right)\left[\hat{R}^{+}_{i}\hat{R}^{-}_{j},\hat{\rho}\right]$
@@ III Equilibrium conditions
$F_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{2}\left\{\left[1-\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\frac{\sin\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.+\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\left[\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}-\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (31)
$\Omega_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{4}\left\{\left[1-\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\frac{\cos\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.-\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}\right)^{2}\right]\left[\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}+\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (32)
and where $\hat{\rho}$ is the density matrix for the atomic system, $\Delta r_{ij}=\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|$ is the distance between Atoms $i$ and $j$, $\bm{\epsilon}_{r}=\left(\mathbf{r}_{i}-\mathbf{r}_{j}\right)/\Delta r_{ij}$ and $\Gamma$ is given by equation (18).
For our two-atom problem the elements of the density matrix will be defined using the basis {|$\left\{\Ket{1}\equiv\Ket{1,0}\right.$, |$\Ket{2}\equiv\Ket{0,0}$, |$\left.\Ket{3}\equiv\Ket{1,-1}\right\}$, where |$\Ket{r,m}$ is given in equations (7)-(9) with $\cos\theta=0$. That is, we have $\rho_{ij}=\braket{i}{\hat{\rho}}{j}$. The solution of the master equation for the two-atom problem discussed in Sec. II.1 is presented in Figure 2 for $\Delta r=0.02\lambda$ with the black curves tracing the time evolution of the corresponding populations ($\lambda$ is the radiation wavelength). The initial conditions were set to $\rho_{11}\left(0\right)=\rho_{22}\left(0\right)=1/2$ and all other populations and coherences to zero, as well as $\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{r}=0$.
As can be inferred from the figure, no energy is trapped in the system as the populations corresponding to the |$\Ket{1,0}$ and |$\Ket{0,0}$ states go to zero in the steady-state. The first (fast) state does so quickly (on a time-scale of $\Gamma t\sim 1/2$) while the second (dark) state decays on a much longer time-scale, but will still eventually go down to zero.
The leakage of internal energy is stopped in the steady-state by enforcing equilibrium conditions for the populations. To see how this happens we modify the time evolution equations for the two-atom populations obtained through the master equation by adding a relaxation term acting on a time-scale $T_{1}$ and a constant pump term, which we then define as $\rho_{jj}^{\mathrm{eq}}/T_{1}$, with $\rho_{jj}^{\mathrm{eq}}$ the equilibrium condition. We thus write
@@
In Figure 2 we also show, using the colored curves and legend, the two-atom system response when $\rho_{11}^{\mathrm{eq}}=\rho_{22}^{\mathrm{eq}}=1/2$, $\rho_{33}^{\mathrm{eq}}=0$ and $\Gamma T_{1}=10$. The energy trapping in the two $m=0$ states is readily observed, with $\approx 0.5$ of the total energy remaining in $m=0$ states and a resulting energy trapping efficiency $\eta\simeq 0.46$.
Such analysis can be extended to cases involving a larger number of atoms. It will then generally be the case that the coherences $\rho_{ij}$, with $i\neq j$, will come into play (they do not affect our calculations for the two-atom populations). We therefore follow the usual practice of introducing a dephasing term scaling as $-\rho_{ij}/T_{2}$, with $T_{2}$ the corresponding time-scale, to account for external factors acting to reduce coherence in the system.
Figure 3: The master equation solution for a system consisting of two two-atom pairs located a distance $0.1\lambda$ apart; the two atoms forming one pair are still separated by $\Delta r=0.02\lambda$. Only the first two-atom pair is initially excited, equally in its fast and dark states. We set the equilibrium conditions for the populations of the four $m=-1$ states (i.e., |$\ket{aa}\ket{1,0}$, |$\ket{aa}\ket{0,0}$, etc.) to $1/4$ and that of the ground state to zero. The black curves show the time evolution of the relevant populations for $\Gamma T_{1}=\Gamma T_{2}=10$, while the red and blue curves show that of the ground state for $\Gamma T_{2}=1$ and $0.1$, respectively; $T_{1}$ is unchanged. Although the increased decoherence reduces the energy trapping efficiency from $\approx 0.6$ to $\approx 0.5$, energy trapping still occurs.
[…]
The effect of the dephasing time-scale is also depicted in the figure with the red and blue curves for the population of the |$\ket{aa}\ket{aa}$ state when $\Gamma T_{2}=1$ and $0.1$, respectively; $T_{1}$ is unchanged. It is found, as could be expected, that increased decoherence (i.e., a lower $T_{2}$) reduces the energy trapping efficiency, but it does not eradicate the effect. In fact, as can be inferred from the results, and a comparison with that from the single two-atom pair of Figure 2, decoherence tends to render the two pairs more uncorrelated.
Figure 4: Intensities (normalized to $\hbar\omega_{0}\Gamma$) for the single-atom, and two- and four-atom cases of Figures 2 and 3 when $\Gamma T_{1}=\Gamma T_{2}=10$. While equilibrium conditions have enable the trapping of energy in the long run, it also brings a non-zero steady-state intensity in all cases. Subradiance and energy trapping are apparent from the narrower profiles of the transient intensities for the two- and four-atom systems over the single-atom case.
Figure 4 shows the intensities (normalized to $\hbar\omega_{0}\Gamma$) for the single-atom, and the two- and four-atom cases of Figures 2 and 3 when $\Gamma T_{1}=\Gamma T_{2}=10$. While equilibrium conditions have enabled the trapping of energy in the long run, it also brings a non-zero steady-state intensity in all cases. This is not surprising since, as discussed above, equilibrium forces the establishing of non-zero steady-state populations. The systems then radiate in this regime at intensities proportional to the sum of the products of the populations and their transitions rates (see equation 19). In the figure all three systems radiate at approximately similar steady-states intensities because of the relatively high value of $\Gamma T_{1}=10$; lower values would generally differentiate between them.
There is, however, another way toward efficient energy trapping. As was observed in Sec. II.1, environmental conditions can sometimes favor the realization of specific radiation modes. In such cases, radiation can become coherent and transition rates can approach those of the small sample systems (see equations 11-14 or 24-27). We now consider this possibility.
IV Superabsorption, coherence and subradiance
Extending the analysis of Sec. III would quickly become prohibitive for even a moderate number of atoms. Furthermore, we have limited ourselves to the emission of a single photon from the $m=-n/2+1$ states to the $m=-n/2$ ground state. Realistic conditions would require initiating the system at higher $m$ values. In order to generalize our analysis we now move to the framework of the Maxwell-Bloch equations (MBE).
In this section, we study the propagation of an incident radiation field on a slab composed of an atomic gas, similar to what is often done in elementary radiative transfer analysis in astrophysics [25]. This will allow us to compare the outcome of our analysis with well-known results. The MBE are perfectly suited for this problem, being one-dimensional in nature. We will also investigate effects the incident field will have on the radiation emanating from the gas itself.
Assuming propagation along the $z$-axis, the MBE [20, 6, 13] for a magnetic dipole transition are [7]
@@ IV Superabsorption, coherence and subradiance
where $n^{\prime}=\left(n_{b}-n_{a}\right)/2$ with $n_{a}$ and $n_{b}$ the densities of atoms in the lower and upper states, respectively, and $M^{+}$ and $B^{+}$ the amplitudes of the molecular magnetization and the magnetic induction field. The temporal derivatives are in relation to the retarded time $\tau=t-z/c$. The ‘slowly varying envelope approximation’ is used with the magnetization and magnetic field vectors given by
$\mathbf{M}^{\pm}\left(z,\tau\right)$ $=M^{\pm}\left(z,\tau\right)e^{\pm i\omega_{0}\tau}\bm{\epsilon}_{\mu}$ (42)
$\mathbf{B^{\pm}}\left(z,\tau\right)$ $=B^{\pm}\left(z,\tau\right)e^{\mp i\omega_{0}\tau}\bm{\epsilon}_{\mu}.$ (43)
As in Sec. III, we added relaxation and dephasing terms with corresponding time-scales $T_{1}$ and $T_{2}$. We can relate this framework to that used in previous sections through the correspondences
@@
at a given position $z$ with $\delta\mathcal{V}$ the volume occupied by the gas in the vicinity. In what follows, we will consider $n_{-}\equiv 2n^{\prime}=n_{b}-n_{a}$ instead of $n^{\prime}$. Furthermore, we will mainly be concerned with conditions of thermal equilibrium or, more generally, situations where $n_{b}<n_{a}$. We thus have, unless otherwise indicated, $n_{-}<0$.
With the incident radiation field propagating along the $z$-axis, as expressed in equation (43) and in the definition of the retarded time, we have an example of situations where environmental or external conditions favor one radiation mode over others. Importantly, the interaction between the incident field and the gas will initially elicit superabsorption and favor coherence in subsequent transitions in the gas through angular correlation among photons, as discussed at the end of Sec. II.1. It is therefore expected that the radiation field emanating from, and propagating within, the gas itself will also take the form of equation (43).
@@
Hence, we will subject the system to two independent stimuli: i) the aforementioned incident magnetic field $B_{0}\equiv B^{+}\left(0,\tau\right)$, which effectively serves as a boundary condition at the input $z=0$ of the system, and ii) internal quantum fluctuations throughout the system at $\tau=0$ modeled as an initial magnetization $M^{+}\left(z,0\right)=n_{-}^{\mathrm{eq}}\mu\,\theta_{0}/2$, where $\theta_{0}=\sqrt{2N}/\left|N_{-}^{\mathrm{eq}}\right|$ is the initial “rising” angle [20] with $N=N_{a}+N_{b}$ the total number of atom ($N_{a}$ and $N_{b}$ in the lower and upper states, respectively) and $N_{-}^{\mathrm{eq}}=N_{b}^{\mathrm{eq}}-N_{a}^{\mathrm{eq}}$ (see below and Appendix B). Unlike for our discussion of Dicke’s work in Sec. II, we express numbers of atoms using capitalized letters from now on to avoid potential confusion with number densities.

### v2

@@ A.1 Two-atom transition rates from $m=0$ and superabsorption
with $\theta$ and $\theta^{\prime}$ denoting the orientations of the first and second absorbed photons, respectively, relative to the $z$-axis. Here again, there is angular correlation, but this time between successively absorbed photons.
A.2 Energy trapping efficiency for $n$ atoms
For the two-atom problem only two permutation symmetries are available for the Dicke states: the totally symmetric species is shared by the $r=1$ triplet (equations 1-3), while the lone $r=0$ singlet is anti-symmetric (equation 4). In the more general case where the gas is composed of an arbitrary number of atoms $n$ a larger number of symmetries, tied to the cooperative number $r$, exist for the Dicke states.
Although it is in principle possible to extend the spontaneous emission rates analysis for finite separations performed for the two-atom case to the $n$-atom problem using appropriate Dicke states, the calculations quickly become prohibitive even for a restricted number of radiators. It is, however, possible to determine the ratio of numbers of photons emitted in the small and infinite-size samples when $m=0$ initially, as done in equation (11) for the two-atom problem. For this, we use the known level of degeneracy of states as a function of the cooperative number $r$ [1]
@@ A.2 Energy trapping efficiency for $n$ atoms
with the fact that, for $n$ even, the number of photons emitted from $m=0$ for the small and infinite-size samples are $r$ and $n/2$, respectively. We thus find for the energy trapping efficiency
@@
One can readily verify that $\eta=1/2$ for the two-atom case, while it is $0.69$ for $n=10$, $0.88$ for $n=100$, etc. Furthermore, it is known that $\overline{r\left(r+1\right)}\simeq m^{2}+n/2$ whenever $n\gg 1$ and $\hbar\omega_{0}\ll k_{\mathrm{B}}T_{\mathrm{kin}}$ we have [1], which from equation (42) yields $\eta\approx 1-\sqrt{2/n}$ when $m=0$.
Appendix B The master equation
The master equation used for the analysis presented in Sec. 2.2 is as follows
$\frac{d\hat{\rho}}{dt}$ $=\frac{1}{i\hbar}\hbar\omega_{0}\sum_{j}\left[\hat{R}^{3}_{j},\hat{\rho}\right]-\frac{\Gamma}{i}\sum_{i\neq j}\Omega_{ij}\left(k\Delta r_{ij}\right)\left[\hat{R}^{+}_{i}\hat{R}^{-}_{j},\hat{\rho}\right]$
@@ Appendix B The master equation
$F_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{2}\left\{\left[1-\left(\boldsymbol{\epsilon}_{\mu}\cdot\boldsymbol{\epsilon}_{r}\right)^{2}\right]\frac{\sin\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.+\left[1-3\left(\boldsymbol{\epsilon}_{\mu}\cdot\boldsymbol{\epsilon}_{r}\right)^{2}\right]\left[\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}-\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (44)
$\Omega_{ij}$ $\left(k\Delta r_{ij}\right)=\frac{3}{4}\left\{\left[1-\left(\boldsymbol{\epsilon}_{\mu}\cdot\boldsymbol{\epsilon}_{r}\right)^{2}\right]\frac{\cos\left(k\Delta r_{ij}\right)}{k\Delta r_{ij}}\right.$
$\left.-\left[1-3\left(\boldsymbol{\epsilon}_{\mu}\cdot\boldsymbol{\epsilon}_{r}\right)^{2}\right]\left[\frac{\sin\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{2}}+\frac{\cos\left(k\Delta r_{ij}\right)}{\left(k\Delta r_{ij}\right)^{3}}\right]\right\}$ (45)
and where $\hat{\rho}$ is the density matrix for the atomic system, $\Delta r_{ij}=\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|$ is the distance between atoms $i$ and $j$, $\boldsymbol{\epsilon}_{r}=\left(\mathbf{r}_{i}-\mathbf{r}_{j}\right)/\Delta r_{ij}$ and $\Gamma$ is given by equation (34). In all calculations presented in Sec. 2.2 we set $\boldsymbol{\epsilon}_{\mu}\cdot\boldsymbol{\epsilon}_{r}=0$ for simplicity.
The time evolution equations for the two-atom populations presented in Figure 2 (black curves) are obtained with equations (43)-(45). Equilibrium conditions are imposed by adding a relaxation term acting on a time-scale $T_{1}$ and a constant pump term, which we then define as $\rho_{jj}^{\mathrm{eq}}/T_{1}$, with $\rho_{jj}^{\mathrm{eq}}$ the equilibrium condition for the corresponding population. We thus write
@@
Appendix C The Maxwell-Bloch equations
Assuming propagation along the $z$-axis, the Maxwell-Bloch equations (MBE) [20, 5, 12] for a magnetic dipole transition are [6]
@@ Appendix C The Maxwell-Bloch equations
where $n^{\prime}=\left(n_{b}-n_{a}\right)/2$ with $n_{a}$ and $n_{b}$ the densities of atoms in the lower and upper states, respectively, and $M^{+}$ and $B^{+}$ the amplitudes of the atomic magnetization and the magnetic induction field. The temporal derivatives are in relation to the retarded time $\tau=t-z/c$. The “slowly varying envelope approximation” is used with the magnetization and magnetic field vectors given by
$\mathbf{M}^{\pm}\left(z,\tau\right)$ $=M^{\pm}\left(z,\tau\right)e^{\pm i\omega_{0}\tau}\boldsymbol{\epsilon}_{\mu}$ (55)
$\mathbf{B^{\pm}}\left(z,\tau\right)$ $=B^{\pm}\left(z,\tau\right)e^{\mp i\omega_{0}\tau}\boldsymbol{\epsilon}_{\mu}.$ (56)
As for the master equation (see equations 46-48), we added relaxation and dephasing terms with corresponding time-scales $T_{1}$ and $T_{2}$, as well as enforcing equilibrium conditions through $n^{\prime}_{\mathrm{eq}}$. We can relate this framework to that used in Secs. 2.1 and 2.2 through the correspondences
@@
at a given position $z$ with $\delta\mathcal{V}$ the volume occupied by the gas in the vicinity. In our analysis and discussion we consider $n_{-}\equiv 2n^{\prime}=n_{b}-n_{a}$ instead of $n^{\prime}$. Furthermore, we are mainly concerned with conditions of thermal equilibrium or, more generally, situations where $n_{b}<n_{a}$. We thus have, unless otherwise indicated, $n_{-}<0$.
@@
As stated in Sec. 2.3, the system is subjected to an incident magnetic field $B_{0}\equiv B^{+}\left(0,\tau\right)$, which effectively serves as a boundary condition at the input $z=0$ of the system, and internal quantum fluctuations throughout the system at $\tau=0$. These fluctuations are modeled as an initial magnetization $M^{+}\left(z,0\right)=n_{-}^{\mathrm{eq}}\mu\,\theta_{0}/2$, where $\theta_{0}=\sqrt{2N}/\left|N_{-}^{\mathrm{eq}}\right|$ is the initial “rising” angle [20] with $N=N_{a}+N_{b}$ the total number of atoms ($N_{a}$ and $N_{b}$ in the lower and upper states, respectively) and $N_{-}^{\mathrm{eq}}=N_{b}^{\mathrm{eq}}-N_{a}^{\mathrm{eq}}$ (see Appendix C.1). Unlike for our discussion of Dicke’s work in Sec. 2.1 and Appendix A, we express numbers of atoms using capitalized letters starting in Sec. 2.3 to avoid potential confusion with number densities.